A locomotive of 60 tone mass pulls some carriages of total mass 240 tones a distance of half a kilometer up a 1 in 200 incline with an acceleration of 0.2m/s. The overall rolling resistance is 5% of normal reaction. What is the total work done

1
Expert's answer
2022-02-17T04:35:54-0500

"A=Fl=(Ma+(M+m)g(\sin\alpha+\mu\cos\alpha))l=87~MJ."